in order to make 0.100 L of a 6.0 M solution of H2SO4, how many grams of H2SO4 must be weighed out? Assume a molar mass of 98.1 g/mol and give the answer to 2 significant figures.
1
Expert's answer
2019-08-13T03:21:07-0400

Let's found how many moles of H2SO4 are there in 0.1L of 6.0M solution:


"\nu = C * V = 6 * 0.1 = 0.6 (moles)"

Find the mass of 0.6 moles of H2SO4:


"m = \nu * M = 0.6 * 98.1 = 59 (g)"

Answer: 59 g of H2SO4 must be weighed out in order to make 0.100 L of a 6.0 M solution.